# Controlled LLM Training on Spectral Sphere

### v1

@@ 2.1 Maximal Update Parametrization ($\bm{\mu}$P)
yang2023spectral shows that preserving scale-invariant activations for feature learning requires the same law to hold for both weights and their updates: $\|{\bm{W}}\|_{2}=\Theta\left(\sqrt{\frac{d_{\mathrm{out}}}{d_{\mathrm{int}}}}\right),\qquad\|{\bm{\Phi}}\|_{2}=\Theta\left(\sqrt{\frac{d_{\mathrm{out}}}{d_{\mathrm{int}}}}\right).$
@@ 2.2 Steepest Descent under Different Norms
Following metrized deep learning (bernstein2024oldoptimizernewnorm), many successful optimizers (e.g. AdamW (loshchilov2019decoupled), Shampoo (gupta2018shampoo), Prodigy (mishchenko2024prodigy)) can be interpreted as first-order methods without convexity assumptions. Specifically, after switching off exponential moving averages, these algorithms reduce to instances of steepest descent governed by distinct norms. Under this framework, an optimizer is fundamentally defined by a weight update $\bm{\Phi}$} that minimizes a quadratic model of the loss:
${\bm{\Phi}}:=\underset{{\bm{\Phi}}}{\operatorname*{argmin}}\left\{\mathcal{L}({\bm{W}})+\langle{\bm{G}},{\bm{\Phi}}\rangle+\frac{\lambda}{2}\|{\bm{\Phi}}\|^{2}\right\}.$ (3)
The update is thus determined by two priors: a norm $\|\!\cdot\!\|$ assigned according to the specific functional role of the module, and a sharpness $\lambda$ that governs the update scale. The solution is as follows:
Given a norm $\|\!\cdot\!\|$ that endows the parameter space with a geometry, the steepest descent update is ${\bm{\Phi}}=-\eta\cdot{\bm{\Phi}}_{\text{unit}},\quad\text{where}\quad{\bm{\Phi}}_{\text{unit}}:=\underset{\|{\bm{T}}\|\leq 1}{\arg\max}\langle{\bm{G}},{\bm{T}}\rangle,\quad\text{and}\quad\eta:=\frac{\|{\bm{G}}\|_{\dagger}}{\lambda}.$ (4) Here, $\|{\bm{G}}\|_{\dagger}:=\max_{\|{\bm{T}}\|\leq 1}\langle{\bm{G}},{\bm{T}}\rangle$ is the dual norm of the gradient $\bm{G}$} induced by $\|\!\cdot\!\|$.
This perspective unifies diverse algorithms by mapping them to their underlying geometries: SGD corresponds to Frobenius norm $\|\!\cdot\!\|_{F}$, Adam to $l_{\infty}$ norm, and Shampoo to spectral norm $\|\!\cdot\!\|_{2}$.
@@ 3.2 First-Order Tangent Space Constraint
$\max_{{\bm{\Phi}}}\ \langle{\bm{G}},{\bm{\Phi}}\rangle\quad\text{s.t.}\quad\|{\bm{\Phi}}\|_{2}\ =1,\ \ \langle{\bm{\Theta}},{\bm{\Phi}}\rangle=0.$ (10)
We then introduce a Lagrange multiplier $\lambda$ and maximize Lagrangian $\mathcal{L}({\bm{\Phi}},\lambda)=\langle{\bm{G}}+\lambda{\bm{\Theta}},{\bm{\Phi}}\rangle$ under constraint $\|{\bm{\Phi}}\|_{2}=1$. The analytical solution and numerical method are summarized below.
@@ 3.3 Second-Order Manifold Constraint
The retraction strictly constrains $\|{\bm{W}}\|_{2}=R$, which via Equation˜14, automatically bounds the weight magnitudes. As a result, weight decay, which is primarily introduced to limit the weight scale, becomes redundant. We therefore eliminate weight decay in hidden 2D weights [^4] We still apply weight decay to 1D params (e.g. Embedding, RMSNorm) for possible scaling stability, although our ablations on 1.7B model suggest that disabling 1D params may actually be slightly better., removing a sensitive hyperparameter from training. More details can be found in Section˜A.3.
@@ 4.1 Spectral Radius Scale
$R=\Theta\left(\sqrt{d_{\text{out}}/d_{\text{in}}}\right)=c\left(\sqrt{d_{\text{out}}/d_{\text{in}}}\right)$
@@ 4.2 Learning Rate Scaler
We propose a unified view in which each learning-rate scaler enforces a consistent effective step size under a chosen norm metric and initialization scheme.
@@ 4.3 Module Granularity
${\bm{W}}=c\sqrt{\frac{d_{\mathrm{out}}}{d_{\mathrm{in}}}}\cdot\frac{{\bm{W}}_{k}}{\|{\bm{W}}_{k}\|_{2}},$
@@ 5.3 Future Improvements
  • Low-precision Training. While weight matrics are spectral constrained, we found the residual stream remains the primary source of outliers. We aim to explore “fully manifold constrained” architectures, i.e. using mHC (xie2025mhcmanifoldconstrainedhyperconnections). Additionally, given SSO’s demonstrated stability, we plan to explore low-precision training (e.g. FP8/NVFP4) to leverage the high throughput of low-bit arithmetic for better training efficiency.
@@ 6.1 Experimental Setup
Following the $\bm{\mu}$P protocol, we perform a learning rate sweep on the 1.7B model with AdamW between $10^{-3}$, $10^{-2}$] and found $5\times 10^{-3}$ to be the optimal value. Training uses 500 warmup steps, a global batch size of $\sim$4M tokens (1024 sequences $\times$ 4096 tokens each), and a cosine learning rate decay reduced to $10\%$ peak. We use BF16 mixed-precision training.
@@ 6.3 MoE 8B-A1B
Router and Load Balancing. The router is implemented in FP32 precision. We adopt the auxiliary-loss-free strategy (wang2024auxiliarylossfreeloadbalancingstrategy), which demonstrated superior expert loading balance compared to global auxiliary loss (qiu2025demonsdetailimplementingload) in our ablations. We enable expert bias with an update rate of 0.001. The sequence-level auxiliary loss is removed, as we found it redundant when expert bias is enabled. we use a sigmoid gate with top-$k$ scores renormalized and scaled by 2. (see Section˜A.4). To evaluate router load balancing, we employ Maximal Violation (MaxVio) (wang2024auxiliarylossfreeloadbalancingstrategy), where 0 indicates perfect balance. As shown in Figure˜9, weight spectral normalization effectively promotes balanced routing, leading to superior validation loss.
@@ Acknowledgements
We sincerely thank Jianlin Su for his excellent blog, Scientific Spaces.
@@ A.5 $\bm{\mu}$P Width Scaling
In Figure˜2, we conduct experiment to validate $\bm{\mu}$P width scaling across different model sizes from 70M to 1.8B. Additional AdamW results are included in Figure˜12. We scale the hidden size, intermediate size, and number of attention heads (while fixing head dimensions). The models are trained on 30B tokens sampled from the Olmo2-mix124 dataset [olmo20252olmo2furious]. All optimizers use the Spectral $\bm{\mu}$P LR Scaler.

### v2

@@ 2.1 Maximal Update Parametrization ($\bm{\mu}$P)
yang2023spectral shows that preserving scale-invariant activations for feature learning requires the same law to hold for both weights and their updates: $\|{\bm{W}}\|_{2}=\Theta\left(\sqrt{\frac{d_{\mathrm{out}}}{d_{\mathrm{in}}}}\right),\qquad\|{\bm{\Phi}}\|_{2}=\Theta\left(\sqrt{\frac{d_{\mathrm{out}}}{d_{\mathrm{in}}}}\right).$
@@ 2.2 Steepest Descent under Different Norms
Following bernstein2024oldoptimizernewnorm; cesista2025sdcbs, many successful optimizers can be interpreted as first-order methods without convexity assumptions. Specifically, after switching off exponential moving averages, these algorithms reduce to instances of steepest descent governed by distinct norms: SGD corresponds to the Frobenius norm $\|\!\cdot\!\|_{F}$, AdamW (loshchilov2019decoupled) to the $\ell_{\infty}$ norm, and Shampoo (gupta2018shampoo) to the Spectral norm $\|\!\cdot\!\|_{2}$.
In this view, we determine the update $\bm{\Phi}$} by iteratively minimizing a first-order Taylor approximation of the loss around the current weights $\bm{W}$}, subject to a hard constraint on the step size:
${\bm{\Phi}}:=\underset{\|{\bm{\Phi}}\|\leq\eta}{\operatorname*{argmin}}\left\{\mathcal{L}({\bm{W}})+\left\langle\nabla_{\bm{W}}\mathcal{L}({\bm{W}}),{\bm{\Phi}}\right\rangle\right\}=\underset{\|{\bm{\Phi}}\|\leq\eta}{\operatorname*{argmin}}\langle{\bm{G}},{\bm{\Phi}}\rangle,$ (3)
where ${\bm{G}}:=\nabla_{\bm{W}}\mathcal{L}({\bm{W}})$ is the gradient of the loss function at $\bm{W}$}. The update is thus determined by two priors: a norm $\|\!\cdot\!\|$ assigned according to the specific functional role of the module, and a step size $\eta$ that governs the update scale. The solution is as follows:
Given a norm $\|\!\cdot\!\|$ defining the geometry and a step size $\eta$ governing the scale, the steepest descent update solution to Equation˜3 is given by ${\bm{\Phi}}=-\eta\cdot{\bm{\Phi}}_{\text{unit}},\quad\text{where}\quad{\bm{\Phi}}_{\text{unit}}:=\underset{\|{\bm{T}}\|\leq 1}{\arg\max}\langle{\bm{G}},{\bm{T}}\rangle,$ (4)
@@ 3.2 First-Order Tangent Space Constraint
$\max_{{\bm{\Phi}}}\ \langle{\bm{G}},{\bm{\Phi}}\rangle\quad\text{s.t.}\quad\|{\bm{\Phi}}\|_{2}\ =1,\ \ \langle{\bm{\Theta}},{\bm{\Phi}}\rangle=0.eq$ (10)
We then introduce a Lagrange multiplier $\lambda$ and maximize Lagrangian $\mathcal{L}({\bm{\Phi}};\lambda)=\langle{\bm{G}}+\lambda{\bm{\Theta}},{\bm{\Phi}}\rangle$ under constraint $\|{\bm{\Phi}}\|_{2}=1$. The analytical solution and numerical method are summarized below.
@@ 3.3 Second-Order Manifold Constraint
The retraction strictly constrains $\|{\bm{W}}\|_{2}=R$, which via Equation˜14, automatically bounds the weight magnitudes. As a result, weight decay, which is primarily introduced to limit the weight scale, becomes redundant. We therefore eliminate weight decay in hidden 2D weights [^4] We still apply weight decay to params like Embedding for possible scaling stability, although our ablations on 1.7B model suggest that disabling weight decay may actually be slightly better., removing a sensitive hyperparameter from training. More details can be found in Section˜A.3.
@@ 4.1 Spectral Radius Scale
$R=\Theta\left(\sqrt{d_{\text{out}}/d_{\text{in}}}\right)=c\left(\sqrt{d_{\text{out}}/d_{\text{in}}}\right),$
@@ 4.2 Learning Rate Scaler
We propose a unified view in which each learning rate scaler enforces a consistent effective step size under a chosen norm metric and initialization scheme.
@@ 4.3 Module Granularity
${\bm{W}}=c\sqrt{\frac{d_{\mathrm{out}}}{d_{\mathrm{in}}}}\cdot\frac{{\bm{W}}_{k}}{\|{\bm{W}}_{k}\|_{2}}.$
@@ 5.3 Future Improvements
  • Low-precision Training. While weight matrics are spectral constrained, we find the residual stream remains the primary source of outliers. We aim to explore “fully manifold constrained” architectures, i.e. using mHC (xie2025mhcmanifoldconstrainedhyperconnections). Additionally, given SSO’s demonstrated stability, we plan to explore low-precision training (e.g. FP8/NVFP4) to leverage the high throughput of low-bit arithmetic for better training efficiency.
@@ 6.1 Experimental Setup
Following the $\bm{\mu}$P protocol, we perform a learning rate sweep on the 1.7B model with AdamW between $10^{-3}$, $10^{-2}$] and find $5\times 10^{-3}$ to be the optimal value. Training uses 500 warmup steps, a global batch size of $\sim$4M tokens (1024 sequences $\times$ 4096 tokens each), and a cosine learning rate decay reduced to $10\%$ peak. We use BF16 mixed-precision training.
@@ 6.3 MoE 8B-A1B
Router and Load Balancing. The router is implemented in FP32 precision. We adopt the auxiliary-loss-free strategy (wang2024auxiliarylossfreeloadbalancingstrategy), which demonstrated superior expert load balancing compared to global auxiliary loss (qiu2025demonsdetailimplementingload) in our ablations. We enable expert bias with an update rate of 0.001. The sequence-level auxiliary loss is removed, as we find it redundant when expert bias is enabled. we use a sigmoid gate with top-$k$ scores renormalized and scaled by 2. (see Section˜A.4). To evaluate router load balancing, we employ Maximal Violation (MaxVio) (wang2024auxiliarylossfreeloadbalancingstrategy), where 0 indicates perfect balance. As shown in Figure˜9, weight spectral normalization effectively promotes balanced routing, leading to superior validation loss.
@@ Acknowledgements
We sincerely thank Jianlin Su for his excellent blog, Scientific Spaces. We also thank Yifei Shen for his thoughtful discussions, and Franz Louis Cesista for helping clarify several conceptual formulations.
@@ A.5 $\bm{\mu}$P Width Scaling
In Figure˜2, we conduct experiment to validate $\bm{\mu}$P width scaling across different model sizes from 70M to 1.8B. Additional AdamW results are included in Figure˜12. We scale the hidden size, intermediate size, and number of attention heads (while fixing head dimensions). The models are trained on 30B tokens sampled from the Olmo2-mix-1124 dataset [olmo20252olmo2furious]. All optimizers use the Spectral $\bm{\mu}$P LR Scaler.